OpenStudy (anonymous):

To find the distance AB across a river, a distance BC=225 is laid off on one side of the river. It is found that B=97 degrees and C=30 degrees. Find AB.

OpenStudy (anonymous):

 \[tg30º=\frac{ x }{ 225 }=\frac{\sqrt{3}}{3}\] \[\frac{x}{75}=\sqrt{3}\]\[x=75\sqrt{3}\]